OpenStudy (anonymous):

prove that \[\cos^4θ−\sin^4θ=\cos^2θ−\sin^2θ\]

OpenStudy (darthsid):

Think about the following two identities: \[a^2 - b^2 = (a+b) (a-b)\] and \[\cos^2\theta + \sin^2\theta = 1\]

OpenStudy (anonymous):

Difference of squares. So, you have: (cos^2(theta) - sin^2(theta))(cos^2(theta) + sin^2(theta)) Since that latter factor is 1 by trig identity, you got the first factor as the result.

OpenStudy (anonymous):

Um Ok, and how would you show that?

OpenStudy (anonymous):

How would you show sin^ + cos^ = 1?

OpenStudy (anonymous):

It's just a rule, isn;t it? I mean, how would you show the workings? Not sure I completely get this topic yet.

OpenStudy (anonymous):

Yes. It is a rule. But, the proof has to do with the right triangle and pythagorean theorem.

OpenStudy (anonymous):

Yes, that's true. But this question didn't want me to answer with the triangle shown... so In algebraic terms, w/o the diagram, how would the steps be shown?

OpenStudy (anonymous):

It is as I showed: (cos^2(theta) - sin^2(theta))(cos^2(theta) + sin^2(theta)) It is like using (a^2-b^2) = (a-b)(a+b) except that here a and b are "sin^2" and "cos^2".

OpenStudy (darthsid):

What GT said is a good way to go, substitute Sin and Cos with 'a' and 'b' to show the working. If you want to prove the rule, here's how Take a right triangle. Let one of the acute angles be A. Let the base (smaller side of the angle) be 'b' Let the perpendicular (side opposite to the angle) be 'p' and Let the hypotenuse (longest side of the triangle) be 'h' Sin A = p/h Cos A = b/h \[\sin^2A = \frac{p^2}{h^2}\] \[\cos ^2A = \frac{b^2}{h^2}\] \[\sin^2A + \cos^2A= \frac{p^2}{h^2} + \frac{b^2}{h^2}\] \[\sin^2A + \cos^2A= \frac{p^2+b^2}{h^2}\] now p^2 + b^2 = h^2 according to pythagoras theorem

OpenStudy (anonymous):

\[Cos^2\theta-\sin^2\theta(\cos^2\theta+\sin^\theta)? \]or \[(Cos^2\theta-\sin^2\theta)(\cos^2\theta+\sin^\theta) ?\]

OpenStudy (anonymous):

latter

OpenStudy (anonymous):

So how (when expanding the bracket) does it simplify to cos^2(theta)-sin^2(theta)?

OpenStudy (anonymous):

Never mind. stupid question

OpenStudy (anonymous):

Thanks for helping! I understand now

OpenStudy (darthsid):

Haha, there are no stupid questions, please feel free to ask.

OpenStudy (anonymous):

Good job!

OpenStudy (anonymous):

I just realized! I would not have thought of it though... need to think more outside the box!

OpenStudy (anonymous):

Sorry again, but again, It wasn't a stupid question... How does it simplify to that answer?

OpenStudy (anonymous):

That is good. Discussing stimulates thought. Also, sometimes not doubting your own good capability helps too.

OpenStudy (anonymous):

I understand the (a-b)(a+b) rule, but how does that simplify it to the second answer... wouldn't it just come back to the first answer?

OpenStudy (darthsid):

So you realize that \[\cos^4\theta - \sin^2\theta = (\cos^2\theta + \sin^2\theta)(\cos^2\theta-\sin^2\theta)\] But since \[\cos^2\theta + \sin^2\theta = 1\] we get \[\cos^4\theta - \sin^2\theta = (1)(\cos^2\theta-\sin^2\theta)\]

OpenStudy (darthsid):

er typo... i meant sin^4 theta in the left side

OpenStudy (anonymous):

Not completely no. I realize the first and second identity, but I'm not sure how it links up!? to the last equation?

OpenStudy (darthsid):

You take the first equation, and put the value of cos^2 + sin^2 from the second equation in it.

OpenStudy (anonymous):

Ohhhhh!

OpenStudy (anonymous):

Thanks. Registration has finally hit!

OpenStudy (darthsid):

That's great! I'm glad I was able to help